# A Review of Barren Plateaus in Variational Quantum Computing

### v1

@@ X Link with vanishing gradient problem
One of the first questions one raises when studying BP phenomena is the relationship to vanishing gradient problems in classical deep learning. Identifying a precise connection and closest analogy between these vanishing gradients has been challenging, in part because there are many different types of vanishing gradients classically. Here we review some of the potential similarities and differences that may lead to a multitude of perspectives.
Differences between the two phenomena. The most well-known variety of classical vanishing gradients is likely the vanishing or exploding norm with depth, caused by growing or shrinking weights due to the activation functions [252]. This phenomenon is already conceptually different than that of BPs in variational quantum computing as (non-noise-induced) BPs arise as the number of qubits increases, not as the depth increases. In fact, for fixed number of qubits there comes a point where adding more layers does not worsen the concentration phenomenon [37, 38].
@@
Some deep classical neural networks still suffer problems even with the solutions listed above. In particular, formally infinite depth networks, such as recurrent neural networks used in language or speech processing, can suffer related problems [256]. Some solutions to this problem are skip-connections, residual connections, and general LSTM models [257]. This strategies could be emulated using simultaneous preparations of quantum states, but direct copying is challenging due to the no-cloning theorem.
Finally, restricted angle range initialization strategies have been shown to lead to larger initialization gradients and can be implemented in both classical and quantum contexts [61, 63, 62, 64, 175]. However, given that quantum landscapes have been shown to have numerous local minima in under-parameterized regimes [16] this strategy might be seen as less promising in quantum contexts.
@@ XI Implications and outlook
The concerted effort by the community has led to significant advancements in the understanding of the BP phenomenon. The primary causes of the phenomenon, which in essence all stem from a curse of dimensionality, have been pinned down. Similarly, progress has been made in identifying architectures that circumvent BPs. However, the ultimate impact of the BP phenomenon on the scalability of variational quantum models is still to be determined.
The questions that remain largely boil down to the fact that the BP phenomenon is an average case notion. If a landscape has a BP then the probability that any randomly chosen parameter setting has non-negligible loss gradients is exponentially small. Given that quantum loss functions are precision limited, i.e., they are estimated from a finite number of measurement shots, the probability of obtaining an informative signal at any randomly chosen parameter setting is exponentially small. Nonetheless, in exponentially narrow regions, a BP landscape can exhibit substantial gradients. The question is whether these regions are easy to find, and even if we find them whether they are useful and can allow for meaningful training. Addressing this will require understanding the complex interplay between the BPs, local minima, expressivity limitations and, most importantly, classical simulability.
At this point, we are thus faced with the question of what next? Classical machine learning has been driven by heuristics and is substantially more successful than what would be expected from theoretical analysis. However, we do not yet have the quantum hardware needed to perform large-scale heuristic implementations of variational quantum models. Indeed, the original motivation for studying BPs was to understand limitations that could prevent scaling variational models to non-classically simulable regimes. In this regard, the field of BP has successfully carved out settings where we know a quantum advantage will be unlikely. Still, much work remains to be done. We envision that the BP study will continue to serve as a guide to determine if, and where, a practical variational quantum advantage exists, and to hint towards the development of new variational methods that go beyond the standard procedure of Fig. 1.

### v2

@@ X Link with vanishing gradient problem
A common question when studying BP phenomena is the relationship to vanishing gradient problems in classical deep learning. Identifying a precise connection between these vanishing gradients has been challenging. This is partly because there are several different types of vanishing gradients in classical settings. Here we review some of the potential similarities and differences that may lead to a multitude of perspectives.
Differences between the two phenomena. The most well-known variety of classical vanishing gradients is the vanishing or exploding norm with depth, caused by growing or shrinking weights due to the activation functions [298]. This phenomenon differs from BPs in variational quantum computing, where (non-noise-induced) BPs occur as the number of qubits increases, not with circuit depth. That is, while quantumly we have a ‘curse of dimension’, classically the problem would more accurately be described as a ‘curse of depth’. In addition, it is worth noting that in the quantum setting, the BP phenomenon can be intimately tied to what $\rho$ and $O$ are, as different initial states and measurements can lead to a model having (or not having) exponentially concentrated loss functions [51, 28].
@@
Some deep classical neural networks still suffer problems even with the solutions listed above. In models like recurrent neural networks, these effects can be examined as the input length increases to infinity, and one can analyze contributions to the current prediction from the infinite past. Unrolling this network leads to a model exhibiting infinite depth, where the problems of vanishing gradients are only further magnified [302]. For a signal to survive from the infinite past, one has to conceive of unitary neural networks that conserve signal at every step, similar to the unitary operations of quantum computers [175]. The general strategy of preserving information from earlier / past layers to improve gradient signals is applicable to networks beyond RNNs, and these solutions include skip-connections, residual connections, and general LSTM models [303]. Strategies derived from the idea of carrying information forward more robustly could be emulated using simultaneous preparations of quantum states, but direct copying is challenging due to the no-cloning theorem.
Finally, restricted angle range initialization strategies [304, 305, 306] and pre-training methods [307] have proven effective classically for avoiding the vanishing gradient problem. Analogous methods have been shown to lead to larger gradients in quantum contexts [68, 70, 69, 71, 202, 204]. However, given that quantum landscapes have been shown to have numerous local minima in under-parameterized regimes [16] this strategy might be seen as less promising in quantum contexts.
@@ XI Implications and outlook
The community’s efforts have led to significant advancements in understanding the BP phenomenon. The primary causes of the phenomenon, which in essence all stem from a curse of dimensionality, have been pinned down. Similarly, progress has been made in identifying architectures that circumvent BPs. However, the ultimate impact of the BP phenomenon on the scalability of variational quantum models is still to be determined.
The questions that remain largely boil down to the fact that the BP phenomenon is an average case notion. If a landscape has a BP then the probability that any randomly chosen parameter setting has non-negligible loss gradients is exponentially small. Given that quantum loss functions are precision limited, i.e., they are estimated from a finite number of measurement shots, the probability of obtaining an informative signal at any randomly chosen parameter setting is exponentially small. Nonetheless, in exponentially narrow regions, a BP landscape can exhibit substantial gradients. The question is whether these regions are easy to find, and even if we find them whether they are useful and can allow for meaningful training. Addressing this requires understanding the interplay between BPs, local minima, expressivity limitations, and, crucially, classical simulability; all of which are open questions that we urge our community to explore. Indeed, we believe that the field of variational quantum computing could benefit from taking a holistic view of all these difficulties and the underlying causes that underpin them.
At this point, we are thus faced with the question of what next? Classical machine learning has been driven by a combination of rigorous theory and well-informed heuristic strategies, leading to successes beyond those predictable by theory alone. However, we do not yet have the quantum hardware needed to perform large-scale heuristic implementations of variational quantum models. Indeed, the original motivation for studying BPs was to understand limitations that could prevent scaling variational models to non-classically simulable regimes. In this regard, the field of BP has successfully carved out settings where we know a quantum advantage will be unlikely. Still, much work remains to be done. We envision that the BP study will continue to serve as a guide to determine if, and where, a practical variational quantum advantage exists, and to hint towards the development of new variational methods that go beyond the standard procedure of Fig. 1.